OpenStudy (anonymous):

g(x)=log(x+5) description of transformation equations for the vertical asymptote(s) x-intercept in (x,y) form

jimthompson5910 (jim_thompson5910):

start with f(x)=log(x) and shift it 5 units to the left to get g(x)=log(x+5) vertical asymptote is x = -5 (since this is the value that makes the argument x+5 equal to zero) x-intercept is (-4,0) and you can find this by plugging in g(x)=0 and solving for x g(x)=log(x+5) 0=log(x+5) 10^1=x+5 1=x+5 1-5=x -4=x x=-4 so again, the x-intercept is (-4,0)

OpenStudy (anonymous):

same thing g(x)=log(-x) description of transformation equation (x,y)

jimthompson5910 (jim_thompson5910):

start with g(x)=log(x) and reflect it over the y-axis to get g(x)=log(-x) vertical asymptote is x=0 x-intercept is (-1,0)

OpenStudy (anonymous):

did you mean start with f(x)=log(x) on the first step

OpenStudy (anonymous):

so there really is no transformation right

jimthompson5910 (jim_thompson5910):

the transformation is the reflection over the y-axis

jimthompson5910 (jim_thompson5910):

yes, you usually start with the parent function, which in this case is f(x)=log(x)

OpenStudy (anonymous):

ok thanks You explained it very well thank you.